In the reaction Mg(OH)2 + 2HCl → MgCl2 + 2H2 O, you were given 15





grams of Mg(OH)2 and 25 grams of HCl. Which of the two reactants is the





limiting reagent? How much MgCl2 will be produced in the reaction? (Mg(OH)2





= 58.32 g/mol, HCl = 36.46 g/mol, MgCl2





= 95.32 g/mol)






1
Expert's answer
2022-03-25T11:25:03-0400

Moles Mg(OH)2: 15 / 58.32 = 0.2572 mol

Moles HCl: 25 / 36.46 = 0.6857 mol

The mole ratio according to the equation is 1 : 2; so, in order for Mg(OH)2 to react completely it requires: 0.2572 * 2 = 0.5144 mol of HCl. We have more (0.6857 mol), so, HCl is the excess reagent, Mg(OH)2 is the limiting one.

Moles of MgCl2 is equal to Mg(OH)2 = 0.2572 mol

Mass = 0.2572 * 95.32 = 24.5 (g)